# Transfer of quantum game strategies

### v1

@@ Abstract.
We develop a method for the transfer of perfect strategies between various classes of two-player, one round cooperative non-local games with quantum inputs and outputs using the simulation paradigm in quantum information theory. We show that such a transfer is possible when canonically associated operator spaces for each game are quantum homomorphic or isomorphic, as defined in [20]. We examine a new class of QNS correlations, needed for the transfer of strategies between games, and characterize them in terms of states on tensor products of canonical operator systems. We define jointly tracial correlations and show they correspond to traces acting on tensor products of canonical ${\rm C}^{*}$-algebras associated with individual game parties. We then make an inquiry into the initial application of such results to the study of concurrent quantum games.
2010 Mathematics Subject Classification. 47L07, 47L25, 81P16, 81R50
@@ 1. Introduction
While our two players are not allowed to communicate during the game, to improve their chances of success they can coordinate their answers according to a predetermined strategy. Our players may have access to a shared quantum entangled state, and measurements on this state by each player can improve their chance of winning by coordinating their answers to the verifier in such a way that would not be possible classically (see [2]: the CHSH inequality and its use in the proof of the celebrated “Bell’s Theorem” during the 1960’s, or [8]). Specific classes of strategies- known as correlations-, based on the use (or non-use) of shared classical and quantum resources between the two players are of particular interest in describing their behavior; such correlations are formalized using classical information channels, when considered as conditional probability distributions. Different mathematical models corresponding to each strategy type describe the outcomes of these experiments: local (corresponding to the use of classical resources), quantum (corresponding to finite-dimensional entangled resources), quantum approximate (corresponding to liminal entangled resources), quantum commuting (which arises from the commuting model of quantum mechanics), and general no-signalling (which does not necessarily rely on the use of a shared resource, but as a probabilistic strategy must still satisfy the basic constraints of the game), are the main correlation classes of interest; these are denoted as $\mathcal{C}_{\rm loc},\mathcal{C}_{\rm q},\mathcal{C}_{\rm qa},\mathcal{C}_{\rm qc}$ and $\mathcal{C}_{\rm ns}$, respectively.
@@
One particular connection between the study of quantum information theory, non-locality, and operator algebras driving much of the recent development in these areas is the equivalence of Tsirelson’s problem in quantum physics, and Connes’ embedding problem (or CEP) in von Neumann algebra theory; this equivalence was established in [17, 22, 30]. The subsequent investigation of this equivalence led to the resolution of many other important questions including a refutation to the weak Tsirelson problem in [34] (see also [15]), and a negative answer to the CEP in [21]. Non-local games lay at the base of all of these approaches, and questions involving non-local games are the main motivation for the current work.
@@
While classical non-local games are fruitful objects of study, as combinatorial objects with finite sets of inputs and outputs they ultimately have inherent restrictions; in an attempt to surpass some of these limitations, more attention has been given to quantum games. These are non-local games where the inputs and outputs are allowed to be quantum states, or sometimes mixtures of classical and quantum states. In this setting, question and answer sets $X,Y,A$ and $B$ are replaced by spaces $\mathbb{C}^{|X|},\mathbb{C}^{|Y|},\mathbb{C}^{|A|}$ and $\mathbb{C}^{|B|}$, and strategies are implemented using quantum channels $\Phi:M_{X}\otimes M_{Y}\rightarrow M_{A}\otimes M_{B}$ instead of classical channels $\mathcal{N}:X\times Y\rightarrow A\times B$. The lack of communication between players is enforced by strictly requiring the use of quantum no-signalling (QNS) channels, as introduced in [14]. Furthermore, the hierarchy of classical correlations is replaced by their quantum analogues, first introduced in [38] (see also [4]). The rules of the game can be generalized to the quantum context by replacing $\lambda$ with a 00-preserving and join-preserving map between the projection lattices $\mathcal{P}_{XY}$ and $\mathcal{P}_{AB}$ of $M_{X}\otimes M_{Y}$ and $M_{A}\otimes M_{B}$, respectively. Such games have been increasingly studied over the past few years (see [6, 7, 11, 33, 38] for a non-exhaustive list). As legitimate generalizations of classical non-local games, the hope is that by “enlarging” the space of possible inputs and outputs (when compared to finite sets), a wealth of new examples are provided that might help shed more light on some of the previously mentioned questions in all areas.
@@
We now describe the organization of the paper in more detail. In Section 2, we set notation and recall the definition of the main no-signalling correlation types for both classical and quantum channels, and introduce the simulation setup for quantum channels. Section 3 contains the definitions of the different types of a class of stochastic (and bistochastic) operator matrices, and collects some of their properties.
These operator matrices will be used to define necessary subclasses of QNS correlations over the quadruple $(X_{2}\times Y_{2})\times(A_{1}\times B_{1})\times(X_{1}\times Y_{1})\times(A_{2}\times B_{2})$ in Section 4, which we call strongly quantum no-signalling; their properties are investigated, leading up to an operator system characterization for these subclasses (similar to those obtained in [28, 38] and [19, Section 7]). In Section 5, the strongly quantum no-signalling correlations are used in the simulation paradigm for quantum channels to establish strategy transport for quantum games; this is achieved in Theorem 5.1 and Theorem 5.2 for the QNS correlation and bicorrelation cases, respectively. We note that the definitions of Sections 3 and 4 generalize those from [19, Section 4, 5]; thus, we recover many of the results (in the classical case) from [19] in Sections 3-5.
All of this setup culminates in Section 6, where we restrict our attention to the transfer of perfect strategies for various types of quantum games. This application employs SQNS correlations as strategies for the ${\rm t}$-homomorphism and isomorphism games between quantum non-local games, when utilizing the framework of generalized homomorphisms of quantum hypergraphs as introduced in [20, Section 3]. A characterization of when perfect strategy transfer is possible between games when the first leg is classical while the second is quantum, along with when both are rank-one quantum games or implication games are included (see [11, 38] for relevant introductions). Furthermore, we also focus on the transfer of strategies between concurrent quantum games- a class of quantum game first introduced in [6] and further developed in [7, 38]. As a proposed quantization of the class of synchronous games, and analogous to the adaptation of the synchronicity condition for classical game homomorphisms which necessitated the definition of jointly synchronous correlations in [19, Section 8], we define jointly tracial SQNS correlations in this section. A tracial characterization of jointly tracial correlations is contained in Theorem 6.16 (in the same vein as the characterizations established in [18], [6] and [7]), and Theorem 6.17 shows that these are the right subclass of SQNS correlations to use for transferring tracial correlations. Thus, the transfer of strategies for quantum games developed within specializes to the concurrent case. Finally, initial connections between the transfer of concurrent/tracial strategies and traces on canonically associated *-algebras and ${\rm C}^{*}$-algebras for concurrent games are discussed.
@@ Acknowledgements
I would like to thank my advisor Ivan G. Todorov for the many fruitful discussions on earlier drafts of the paper. I would also like to thank Lyudmila Turowska for helpful comments on some results in the last section, and Alexandros Chatzinikalaou for identifying the gap in the proof of [19, Lemma 5.8], necessitating the change in definition.
@@ 2. Preliminaries
In this section we set notation, and include the necessary preliminaries on quantum no-signalling correlations to be used throughout the rest of the paper. For a finite set $X$, we let $\mathbb{C}^{X}=\oplus_{x\in X}\mathbb{C}$ and write $(e_{x})_{x\in X}$ for the canonical orthonormal basis of $\mathbb{C}^{X}$. Similarly, if $H$ is a Hilbert space we set $H^{X}=\oplus_{x\in X}H$. If $X$ is countable (not necessarily finite), let $\ell_{2}^{X}$ denote the Hilbert space of square summable sequences over $\mathbb{C}$; in the specific case that $X$ is finite, we will sometimes write $\ell_{2}^{X}=\mathbb{C}^{X}$. For finite $X,Y$, we will often abbreviate the Cartesian product $X\times Y$ as $XY$ for ease of use in notation. We denote by $M_{X}$ the algebra of all complex matrices of size $X\times X$, and by $\mathcal{D}_{X}$ its subalgebra of all diagonal matrices. We write $\epsilon_{xx^{\prime}},x,x^{\prime}\in X$ for the canonical matrix units in $M_{X}$, denote by ${\rm Tr}$ the trace functional on $M_{X}$, and set $\langle S,T\rangle={\rm Tr}(ST^{\rm t})$, where the transpose is with respect to the canonical orthonormal basis. We let $\Delta_{X}:M_{X}\rightarrow\mathcal{D}_{X}$ denote the canonical conditional expectation on the full matrix algebra. Set
@@
and $\tilde{J}_{X}:=\frac{1}{|X|}J_{X}$. If $\mathfrak{m}_{X}=\sum\limits_{x\in X}e_{x}\otimes e_{x}$ is the maximally entangled unit vector in $\mathbb{C}^{X}\otimes\mathbb{C}^{X}$, then $\tilde{J}_{X}=\mathfrak{m}_{X}\mathfrak{m}_{X}^{*}$ is the corresponding rank-one projection. Note that $J_{X}^{\rm cl}=\Delta_{XX}(J_{X})$, and thus we may think of $J_{X}^{\rm cl}$ as the “classical” part of the (unnormalized) state $J_{X}$.
@@
[…]
(1) $\Gamma[\mathcal{E}]=\sum\limits_{i=1}^{k}\Psi_{i}\circ\mathcal{E}\circ\Phi_{i}.$
(2) $\bigg{(}\sum\limits_{x_{1},x_{1}^{\prime}}\sum\limits_{y_{1},y_{1}^{\prime}}\Gamma(x_{1}x_{1}^{\prime},y_{2}y_{2}^{\prime}|x_{2}x_{2}^{\prime},y_{1}y_{1}^{\prime})\mathcal{E}(y_{1}y_{1}^{\prime}|x_{1}x_{1}^{\prime})\bigg{)}_{x_{2},x_{2}^{\prime},y_{2},y_{2}^{\prime}},$
where the internal sums range over $X_{1}$ and $Y_{1}$, for all $x_{2},x_{2}^{\prime}\in X_{2},y_{2},y_{2}^{\prime}\in Y_{2}$.
In the remainder of this section, we recall the basic types of quantum and classical no-signalling correlations that will be used throughout the paper. Let $X,Y,A$ and $B$ be finite sets. A quantum no-signalling (QNS) correlation [14] is a quantum channel $\Gamma:M_{XY}\rightarrow M_{AB}$ such that
(3) ${\rm Tr}_{A}\Gamma(\rho_{X}\otimes\rho_{Y})=0\text{ whenever }\rho_{X}\in M_{X}\text{ and }{\rm Tr}(\rho_{X})=0,$
@@
(4) ${\rm Tr}_{B}\Gamma(\rho_{X}\otimes\rho_{Y})=0\text{ whenever }\rho_{Y}\in M_{Y}\text{ and }{\rm Tr}(\rho_{Y})=0.$
@@
As discussed in [38, Section 3], stochastic operator matrices $E$ are the Choi matrices of unital completely positive maps $\Phi_{E}:M_{A}\rightarrow M_{X}\otimes\mathcal{B}(H)$ given by
(5) $\Phi_{E}(\epsilon_{a,a^{\prime}})=E_{a,a^{\prime}},\;\;\;\;a,a^{\prime}\in A.$
@@
(6) $\Gamma_{E,\sigma}(\rho_{X})=\Phi_{*}(\rho_{X}\otimes\sigma),\;\;\;\;\rho_{X}\in M_{X}.$
@@
for all $x,x^{\prime}\in X,y,y^{\prime}\in Y,a,a^{\prime}\in A,b,b^{\prime}\in B$, we let $E\cdot F$ be the (unique) stochastic operator matrix over $(XY,AB)$ (see [38, Proposition 4.1]) defined by
@@
A QNS correlation $\Gamma:M_{XY}\rightarrow M_{AB}$ is called quantum commuting if there exists a Hilbert space $H$, a unit vector $\xi\in H$ and stochastic operator matrices $\tilde{E}=(E_{x,x^{\prime},a,a^{\prime}})_{x,x^{\prime},a,a^{\prime}}$ and $\tilde{F}=(F_{y,y^{\prime},b,b^{\prime}})_{y,y^{\prime},b,b^{\prime}}$ on $H$ such that $\tilde{E}$ and $\tilde{F}$ are mutually commuting, with $\Gamma=\Gamma_{E,F,\xi}$. Quantum QNS correlations are those for which there exist finite dimensional Hilbert spaces $H_{A},H_{B}$, stochastic operator matrices $E\in M_{XA}\otimes\mathcal{B}(H_{A})$ and $F\in M_{YB}\otimes\mathcal{B}(H_{B})$, and a pure state $\sigma\in\mathcal{T}(H_{A}\otimes H_{B})$ such that $\Gamma=\Gamma_{E\odot F,\sigma}$. Approximately quantum QNS correlations are the limits of quantum QNS correlations, while local QNS correlations are the convex combinations of the form $\Gamma=\sum_{i=1}^{k}\lambda_{i}\Phi_{i}\otimes\Psi_{i}$, where $\Phi_{i}:M_{X}\to M_{A}$ and $\Psi_{i}:M_{Y}\to M_{B}$ are quantum channels, $i=1,\dots,k$. We write $\mathcal{Q}_{\rm qc}$ (resp. $\mathcal{Q}_{\rm qa}$, $\mathcal{Q}_{\rm q}$, $\mathcal{Q}_{\rm loc}$) for the (convex) set of all quantum commuting (resp. approximately quantum, quantum, local) QNS correlations, and note the (strict, see [38]) inclusions
(7) $\mathcal{Q}_{\rm loc}\subseteq\mathcal{Q}_{\rm q}\subseteq\mathcal{Q}_{\rm qa}\subseteq\mathcal{Q}_{\rm qc}\subseteq\mathcal{Q}_{\rm ns}.$
@@
[…]
may be considered as an operator subsystem of $M_{XX}$. By [38, Proposition 5.5, Theorem 6.2], the elements $\Gamma$ of $\mathcal{Q}_{\rm ns}$ correspond canonically to elements of the tensor product $\mathcal{L}_{X,A}\otimes_{\min}\mathcal{L}_{Y,B}$ (viewed as an operator subsystem of $M_{XA}\otimes M_{YB}$), and elements in $\mathcal{Q}_{\rm ns}^{\rm bi}$ corresponding to elements of $\mathcal{L}_{X}\otimes_{\rm min}\mathcal{L}_{Y}$ [7, Proposition 3.6, Theorem 5.4].
@@
(8) $\mathcal{N}_{p}(\rho)=\sum\limits_{x\in X,y\in Y}\sum\limits_{a\in A,b\in B}p(a,b|x,y)\langle\rho(e_{x}\otimes e_{y}),e_{x}\otimes e_{y}\rangle\epsilon_{aa}\otimes\epsilon_{bb}.$

### v2

@@ Abstract.
[…]
Key words and phrases:
Non-local game, quantum no-signalling correlations, operator systems, game algebra.
2020 Mathematics Subject Classification:
47L07, 47L25, 81P16, 81R50
@@ 1. Introduction
While our two players are not allowed to communicate during the game, to improve their chances of success they can coordinate their answers according to a predetermined strategy. Our players may have access to a shared quantum entangled state, and measurements on this state by each player can improve their chance of winning by coordinating their answers to the referee in such a way that would not be possible classically (see [2]: the CHSH inequality and its use in the proof of the celebrated “Bell’s Theorem” during the 1960’s, or [8]). Specific classes of strategies— known as correlations— based on the use (or non-use) of shared classical and quantum resources between the two players are of particular interest in describing their behavior; such correlations are formalized using classical information channels, when considered as conditional probability distributions. Different mathematical models corresponding to each strategy type describe the outcomes of these experiments: local (corresponding to the use of classical resources), quantum (corresponding to finite-dimensional entangled resources), quantum approximate (corresponding to liminal entangled resources), quantum commuting (which arises from the commuting model of quantum mechanics), and general no-signalling (which does not necessarily rely on the use of a shared resource, but as a probabilistic strategy must still satisfy the basic constraints of the game), are the main correlation classes of interest. These are denoted as $\mathcal{C}_{\rm loc},\mathcal{C}_{\rm q},\mathcal{C}_{\rm qa},\mathcal{C}_{\rm qc}$ and $\mathcal{C}_{\rm ns}$, respectively.
@@
One particular connection between the study of quantum information theory, non-locality, and operator algebras driving much of the recent development in these areas is the equivalence of Tsirelson’s problem in quantum physics, and Connes’ embedding problem (or CEP) in von Neumann algebra theory; this equivalence was established in [18, 23, 31]. The subsequent investigation of this equivalence led to the resolution of many other important questions including a refutation to the strong Tsirelson problem in [35] (see also [16]), and a negative answer to the CEP in [22]. Non-local games lay at the base of all of these approaches, and questions involving non-local games are the main motivation for the current work.
@@
While classical non-local games are fruitful objects of study, as combinatorial objects with finite sets of inputs and outputs they ultimately have inherent restrictions; in an attempt to surpass some of these limitations, more attention has been given to quantum games. These are non-local games where the inputs and outputs are allowed to be quantum states, or sometimes mixtures of classical and quantum states. In this setting, question and answer sets $X,Y,A$ and $B$ are replaced by spaces $\mathbb{C}^{|X|},\mathbb{C}^{|Y|},\mathbb{C}^{|A|}$ and $\mathbb{C}^{|B|}$, and strategies are implemented using quantum channels $\Phi:M_{X}\otimes M_{Y}\rightarrow M_{A}\otimes M_{B}$ instead of classical channels $\mathcal{N}:X\times Y\rightarrow A\times B$. The lack of communication between players is enforced by strictly requiring the use of quantum no-signalling (QNS) channels, as introduced in [15]. Furthermore, the hierarchy of classical correlations is replaced by their quantum analogues, first introduced in [39] (see also [4]). The rules of the game can be generalized to the quantum context by replacing $\lambda$ with a 0-preserving and join-preserving map between the projection lattices $\mathcal{P}_{XY}$ and $\mathcal{P}_{AB}$ of $M_{X}\otimes M_{Y}$ and $M_{A}\otimes M_{B}$, respectively. Such games have been increasingly studied over the past few years (see [6, 7, 11, 34, 39] for a non-exhaustive list). As legitimate generalizations of classical non-local games, the hope is that by “enlarging” the space of possible inputs and outputs (when compared to finite sets), a wealth of new examples are provided that might help shed more light on some of the previously mentioned questions in all areas.
@@
We now describe the organization of the paper in more detail. In Section 2, we set notation and recall the definition of the main no-signalling correlation types for both classical and quantum channels, and introduce the simulation setup for quantum channels. Section 3 contains the definitions of the different types of a class of stochastic (and bistochastic) operator matrices, and how these operator matrices will be used to define necessary subclasses of QNS correlations (which we call strongly quantum no-signalling) over the quadruple $(X_{2}\times Y_{2})\times(A_{1}\times B_{1})\times(X_{1}\times Y_{1})\times(A_{2}\times B_{2})$. In Section 4, the strongly quantum no-signalling correlations are used in the simulation paradigm for quantum channels to establish strategy transport for quantum games; this is achieved in Theorem 4.1 and Theorem 4.2 for the QNS correlation and bicorrelation cases, respectively. We note that the definitions of Sections 3 and 3.1 generalize those from [20, Section 4, 5]; thus, we recover many of the results (in the classical case) from [20] in Sections 3-4.
Section 5 contains some of the main results of the paper, where we restrict our attention to the transfer of perfect strategies for various types of quantum games. This application employs SQNS correlations as strategies for the $\rm t$}-homomorphism and isomorphism games between quantum non-local games, when utilizing the framework of generalized homomorphisms of quantum hypergraphs as introduced in [21, Section 3]. A characterization of when perfect strategy transfer is possible between games when the first leg is classical while the second is quantum, along with when both are implication games are included (see [39] for relevant introductions).
[…]
@@ Acknowledgements
I would like to thank my advisor Ivan G. Todorov for the many fruitful discussions on earlier drafts of the paper. I would also like to thank Lyudmila Turowska for helpful comments on some results in the last section, and Alexandros Chatzinikolaou for identifying the gap in the proof of [20, Lemma 5.8], necessitating the change in definition. Finally, I would like to thank the referee(s) for their detailed reading and remarks.
@@ 2. Preliminaries
In this section we set notation, and include the necessary preliminaries on quantum no-signalling correlations to be used throughout the rest of the paper. For a finite set $X$, we let $\mathbb{C}^{X}=\oplus_{x\in X}\mathbb{C}$ and write $(e_{x})_{x\in X}$ for the canonical orthonormal basis of $\mathbb{C}^{X}$. Similarly, if $H$ is a Hilbert space we set $H^{X}=\oplus_{x\in X}H$. If $X$ is countable (not necessarily finite), let $\ell_{2}^{X}$ denote the Hilbert space of square summable sequences over $\mathbb{C}$; in the specific case that $X$ is finite, we will sometimes write $\ell_{2}^{X}=\mathbb{C}^{X}$. For finite $X,Y$, we will often abbreviate the Cartesian product $X\times Y$ as $XY$ for ease of use in notation. We denote by $M_{X}$ the algebra of all complex matrices of size $X\times X$, and by $\mathcal{D}_{X}$ its subalgebra of all diagonal matrices. We write $\epsilon_{xx^{\prime}},x,x^{\prime}\in X$ for the canonical matrix units in $M_{X}$, denote by $\rm Tr$} the trace functional on $M_{X}$, and set $\langle S,T\rangle={\rm Tr}(ST^{*})$, where the adjoint is with respect to the canonical orthonormal basis. We let $\Delta_{X}:M_{X}\rightarrow\mathcal{D}_{X}$ denote the canonical conditional expectation on the full matrix algebra. Set
@@
and $\tilde{J}_{X}:=\frac{1}{|X|}J_{X}$. If $\mathfrak{m}_{X}=\frac{1}{\sqrt{|X|}}\sum\limits_{x\in X}e_{x}\otimes e_{x}$ is the maximally entangled unit vector in $\mathbb{C}^{X}\otimes\mathbb{C}^{X}$, then $\tilde{J}_{X}=\mathfrak{m}_{X}\mathfrak{m}_{X}^{*}$ is the corresponding rank-one projection. Note that $J_{X}^{\rm cl}=\Delta_{XX}(J_{X})$, and thus we may think of $J_{X}^{\rm cl}$ as the “classical” part of the (unnormalized) state $J_{X}$.
@@
Let $X$ and $A$ be finite sets. A $\it classical\;information\;channel$} from $X$ to $A$ is a positive trace preserving linear map $\mathcal{N}:\mathcal{D}_{X}\rightarrow\mathcal{D}_{A}$. If $\mathcal{N}$ is an information channel, setting $p(\cdot|x)=\mathcal{N}(\epsilon_{xx})$ for each $x\in X$ it is easy to see that $\mathcal{N}$ is completely determined by its corresponding family of conditional probability distributions $\{(p(a|x))_{a\in A}:\;x\in X\}$.
A $\it quantum\;information\;channel$} from $X$ to $A$ is a completely positive trace preserving linear map $\Phi:M_{X}\rightarrow M_{A}$. A quantum channel will be called $(X,A)$-$\it classical$} if $\Phi=\Delta_{A}\circ\Phi\circ\Delta_{X}$. Any classical channel $\mathcal{N}:\mathcal{D}_{X}\rightarrow\mathcal{D}_{A}$, has a corresponding $(X,A)$-classical (quantum) channel $\Phi_{\mathcal{N}}:M_{X}\rightarrow M_{A}$ given by $\Phi_{\mathcal{N}}=\mathcal{N}\circ\Delta_{X}$. Conversely, any quantum channel $\Phi:M_{X}\rightarrow M_{A}$ induces a classical channel $\mathcal{N}_{\Phi}:\mathcal{D}_{X}\rightarrow\mathcal{D}_{A}$ given by $\Delta_{A}\circ\Phi|_{\mathcal{D}_{X}}$. Finally, if $\mathcal{E}:\mathcal{D}_{X}\rightarrow M_{A}$ is a (classical-to-quantum) channel, set $\Gamma_{\mathcal{E}}=\mathcal{E}\circ\Delta_{X}$, so $\Gamma_{\mathcal{E}}$ is a quantum channel from $M_{X}$ to $M_{A}$.
In the remainder of this section, we recall the basic types of quantum and classical no-signalling correlations that will be used throughout the paper, along with establishing the simulation paradigm arising from quantum information theory. Let $X,Y,A$ and $B$ be finite sets. A quantum no-signalling (QNS) correlation [15] is a quantum channel $\Gamma:M_{XY}\rightarrow M_{AB}$ such that
${\rm Tr}_{A}\Gamma(\rho_{X}\otimes\rho_{Y})=0\text{ whenever }\rho_{X}\in M_{X}\text{ and }{\rm Tr}(\rho_{X})=0,$ (1)
@@
${\rm Tr}_{B}\Gamma(\rho_{X}\otimes\rho_{Y})=0\text{ whenever }\rho_{Y}\in M_{Y}\text{ and }{\rm Tr}(\rho_{Y})=0.$ (2)
@@
As discussed in [39, Section 3], stochastic operator matrices $E$ are the Choi matrices of unital completely positive maps $\Phi_{E}:M_{A}\rightarrow M_{X}\otimes\mathcal{B}(H)$ given by
$\Phi_{E}(\epsilon_{a,a^{\prime}})=E_{a,a^{\prime}},\;\;\;\;a,a^{\prime}\in A.$ (3)
@@
$\Gamma_{E,\sigma}(\rho_{X})=\Phi_{*}(\rho_{X}\otimes\sigma),\;\;\;\;\rho_{X}\in M_{X}.$ (4)
@@
for all $x,x^{\prime}\in X,y,y^{\prime}\in Y,a,a^{\prime}\in A,b,b^{\prime}\in B$, we let $E\cdot F$ be the (unique) stochastic operator matrix over $(XY,AB)$ (see [39, Proposition 4.1]) defined by
@@
A QNS correlation $\Gamma:M_{XY}\rightarrow M_{AB}$ is called quantum commuting if there exists a Hilbert space $H$, a unit vector $\xi\in H$ and stochastic operator matrices $E=(E_{x,x^{\prime},a,a^{\prime}})_{x,x^{\prime},a,a^{\prime}}$ and $F=(F_{y,y^{\prime},b,b^{\prime}})_{y,y^{\prime},b,b^{\prime}}$ on $H$ such that $E$ and $F$ are mutually commuting, with $\Gamma=\Gamma_{E,F,\xi}$. Quantum QNS correlations are those for which there exist finite dimensional Hilbert spaces $H_{A},H_{B}$, stochastic operator matrices $E\in M_{XA}\otimes\mathcal{B}(H_{A})$ and $F\in M_{YB}\otimes\mathcal{B}(H_{B})$, and a pure state $\sigma\in\mathcal{T}(H_{A}\otimes H_{B})$ such that $\Gamma=\Gamma_{E\odot F,\sigma}$. Approximately quantum QNS correlations are the limits of quantum QNS correlations, while local QNS correlations are the convex combinations of the form $\Gamma=\sum_{i=1}^{k}\lambda_{i}\Phi_{i}\otimes\Psi_{i}$, where $\Phi_{i}:M_{X}\to M_{A}$ and $\Psi_{i}:M_{Y}\to M_{B}$ are quantum channels, $i=1,\dots,k$. We write $\mathcal{Q}_{\rm qc}$ (resp. $\mathcal{Q}_{\rm qa}$, $\mathcal{Q}_{\rm q}$, $\mathcal{Q}_{\rm loc}$) for the (convex) set of all quantum commuting (resp. approximately quantum, quantum, local) QNS correlations, and note the (strict, see [39]) inclusions
$\mathcal{Q}_{\rm loc}\subseteq\mathcal{Q}_{\rm q}\subseteq\mathcal{Q}_{\rm qa}\subseteq\mathcal{Q}_{\rm qc}\subseteq\mathcal{Q}_{\rm ns}.$ (5)
@@
 $\text{ and }\sum_{x\in X}\lambda_{x,x,a,a^{\prime}}=\delta_{a,a^{\prime}}c,\;x,x^{\prime},a,a^{\prime}\in X\bigg\}$
may be considered as an operator subsystem of $M_{XX}$. By [39, Proposition 5.5, Theorem 6.2], the elements $\Gamma$ of $\mathcal{Q}_{\rm ns}$ correspond canonically to elements of the tensor product $\mathcal{L}_{X,A}\otimes_{\min}\mathcal{L}_{Y,B}$ (viewed as an operator subsystem of $M_{XA}\otimes M_{YB}$), and elements in $\mathcal{Q}_{\rm ns}^{\rm bi}$ corresponding to elements of $\mathcal{L}_{X}\otimes_{\rm min}\mathcal{L}_{Y}$ [7, Proposition 3.6, Theorem 5.4].
@@
$\mathcal{N}_{p}(\rho)=\sum\limits_{x\in X,y\in Y}\sum\limits_{a\in A,b\in B}p(a,b|x,y)\langle\rho(e_{x}\otimes e_{y}),e_{x}\otimes e_{y}\rangle\epsilon_{aa}\otimes\epsilon_{bb}.$ (6)
@@
$\Gamma[\mathcal{E}]=\sum\limits_{i=1}^{k}\Psi_{i}\circ\mathcal{E}\circ\Phi_{i}.$ (7)
$\bigg(\sum\limits_{x_{1},x_{1}^{\prime}}\sum\limits_{y_{1},y_{1}^{\prime}}\Gamma(x_{1}x_{1}^{\prime},y_{2}y_{2}^{\prime}|x_{2}x_{2}^{\prime},y_{1}y_{1}^{\prime})\mathcal{E}(y_{1}y_{1}^{\prime}|x_{1}x_{1}^{\prime})\bigg)_{x_{2},x_{2}^{\prime},y_{2},y_{2}^{\prime}},$ (8)
where the internal sums range over $X_{1}$ and $Y_{1}$, for all $x_{2},x_{2}^{\prime}\in X_{2},y_{2},y_{2}^{\prime}\in Y_{2}$. Thus, channel simulation is a process for constructing new quantum channels from given ones, with the assistance of correlations; morever, it is natural to want the simulated channel to depend on a shared resource between two parties. If Alice and Bob have access to some shared resource (for instance: shared randomness, or entanglement, etc.), their interaction with the resource yields the no-signalling correlation $\Gamma$, and simulated channel $\Gamma[\mathcal{E}]$ is dependent on their local operations (see [12, Section II], and [15]).